TIME-RESOLVED INFRARED DIODE LASER SPECTROSCOPY OF THE ν_{1} + ν_{2} - ν_{2} HOT BAND OF THE FeCO RADICAL PRODUCED BY THE ULTRAVIOLET LASER PHOTOLYSIS OF Fe(CO)_{5}

\underline{KEIICHI TANAKA}, T. HIKITA, and T. TANAKA, Department of Chemistry, Kyushu University, Fukuoka, 812-8581 Japan.

Time-resolved infrared diode laser spectroscopy was applied to the observation of the ν_{1} + ν_{2} - ν_{2} hot band of the iron carbonyl (FeCO) radical. The FeCO radical was produced by the ArF excimer laser photolysis of iron pentacarbonyl Fe(CO)_{5}. Although the FeCO radical has the ^{3}Σ^{−} electronic ground state, the ν_{2} bending vibrational excited state has vibronically the ^{3}Π symmetry, which allows the vibronic interaction with the ^{3}Π electronic state located very closely, 6400 cm^{−1}, to the ground electronic state ^{a}.

In the frequency region of 1930-1950 cm^{−1}, more than 50 R- and P -branch lines were identified which split into triplet (P = 0, 1, and 2) due to the spin-spin interaction. Each spin component split further into doublet, as Λ-type doubling, because of the vibraonic interaction with the nearby ^{3}Π electronic state. The band origin ν_{0} = 1941.54573 (50) cm^{−1} was determined as well as the molecular constants for the ν_{1} + ν_{2} state, fixing the molecular constants for the ν_{2} state to those of the millimeter-wave ^{b}. results. The effective spin-orbit constant A_{e}ff and the Λ-type doubling constants for the ν_{1} + ν_{2} state are much different from those of the ν_{2} state indicating that the effect of vibronic interaction with the ^{3}Π electronic state has increased due to the excitation of the ν_{1} vibration.

aK. Tanaka, M. Shirasaka, and T. Tanaka, J. Chem. Phys., 106, 6820 (1997).

bK. Tanaka, M. Nakamura, M. Shirasaka, and T. Tanaka, The 53rd Symposium on Molecular Spectroscopy, Columbus Ohio, (1998) MG16.